# Privacy-Preserving Resilient Vector Consensus

### v1

@@ 5.1 $\rho-$Concentrated Geo-Privacy
We analyze the privacy of infinite sequences as outputs for two reasons. First, due to the noise adding, convergence is non-deterministic and can theoretically only be achieved as the iteration approaches infinity, when the noise diminishes to a negligible level. Secondly, since the output at each iteration can potentially leak the initial states to varying degrees, we apply a composition conclusion to analyze the upper bound of privacy leakage. It is important to note that we only consider the states of normal agents and neglect the states of the faulty agents, as we are concerned with the state evolution of normal agents under PP-ADRC. Normal agents may leak privacy when sending state messages; thus, this issue is independent of the status and transmitted information of the faulty agents. However, the states of normal agents can be influenced by the states of faulty agents. Moreover, given variables $\lambda$ and $\upsilon$, we plot the surface of $\rho$ concerning them, as shown in Figure 3. It can be observed that $\rho$ monotonically increases with both $\lambda$ and $\upsilon$.
Figure 3: Change trend of $\rho$
5.2 Comparison with $(\varepsilon,\delta)-$Differential Privacy
We first give the definition of $(\varepsilon,\delta)-$DP in the resilient vector consensus.
@@ 5.2 Comparison with $(\varepsilon,\delta)-$Differential Privacy
Given $\ell\in\mathbb{R}\geq 0$, the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ are $\ell$-neighboring if, for some $i_{0}\in\overline{\mathcal{V}},||[\overline{\mathbf{x}}_{0}]_{i_{0}}-[\overline{\mathbf{x}}_{0}^{\prime}]_{i_{0}}||_{2}\leq\ell$ and $[\overline{\mathbf{x}}_{0}]_{i}=[\overline{\mathbf{x}}_{0}^{\prime}]_{i}$ if $i\neq i_{0}$ and $i\in\overline{\mathcal{V}}$. Given $\ell$, $\varepsilon$ and $\delta\in\mathbb{R}\geq 0$, the system achieves $(\varepsilon,\delta)$-differentially private if, for any pair of $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ of $\ell-$neighboring initial states and any set $\mathcal{O}\in\mathcal{B}((\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}})$,
@@
The definition of neighboring initial states here differs from those in existing papers [21, 20, 29], where the $1$-norm is used. In those papers, differential privacy is $\epsilon$-DP with Laplace noise, whereas we use $(\epsilon,\delta)$-DP with Gaussian noise. Therefore, the neighboring initial states are defined using the $2$-norm.
As we have talked before, the output of the system is an infinite sequence. The output at each time can be considered as a query. Therefore, the final result is the infinite composition of each single query at time $h$ satisfying $(\varepsilon(h),\delta(h))-$DP. We now give the detail of $\varepsilon(h)$ and $\delta(h)$ according to Theorem A.1. in [32].
@@
Regarding the composition, we can simply sum $\varepsilon(h)$ and $\delta(h)$ from $h=0$ to $\infty$. This results in rapid growth of privacy loss and $\delta$, indicating a poor outcome. However, computing the tightest possible privacy guarantee for such a composition is #P-hard [33]. Then, we can make a detailed comparison between these two definitions below:
  • CGP better meets the privacy preservation needs. In resilient vector consensus, our privacy preservation target is the initial states of all normal agents, which may be leaked during communication, rather than privacy leakage caused by the alteration or addition/deletion of an initial state in a data center. Therefore, the definition of $\ell$-neighboring states in DP does not fully meet our needs, as it should consider any pair of initial states. Consequently, CGP describes the degree of output differences for any pair of inputs, making it more suitable for protecting the initial state of each normal agent under resilient vector consensus. It also uses the distance between initial states as a privacy protection parameter, thereby greatly improving the system’s utility.
@@ 5.3 Trade-off between Privacy and Accuracy
In this part, we discuss the tradeoff between privacy and accuracy. As we analyze the convergence accuracy from two perspectives, i.e., distribution of the final value and convex hull change, we also detail the tradeoff from these two perspectives. From Theorem 5.20, the parameters that determine $\rho$ are provided. The parameters that we can freely choose are the noise parameters $\lambda$ and $\upsilon$. The larger $\lambda$ and $\upsilon$ are, the stronger the privacy preservation is. Therefore, we analyze the relationship between accuracy and privacy by changing $\lambda$ and $\upsilon$.
  • Distribution of the final value: The Mahalanobis distance is a normalized coefficient, meaning its value is fixed artificially and remains uninfluenced by variance or covariance. However, the volume of the hyperspheroid formed at an equal Mahalanobis distance can be influenced by various parameters. Therefore, the metric that accurately represents precision should be the volume of the hyperspheroid: the larger the volume, the lower the precision. As previously analyzed, increasing the amplitude parameters $\lambda$ and $\upsilon$ of the noise increases the volume of the hyperspheroid, thereby reducing convergence accuracy and decreasing $\rho$, which corresponds to enhanced privacy protection.
  • Convex hull change: We use the Hausdorff distance to characterize the distance between convex hulls before and after noise adding. Similar to the Mahalanobis distance, the Hausdorff distance is also specified artificially. However, the probability of convergence within the convex hull is influenced by the noise parameters. The larger the noise amplitude parameters $\lambda$ and $\upsilon$, the lower the probability that the new convex hull will converge within a given Hausdorff distance. This results in a decrease in convergence accuracy but simultaneously leads to a reduction in $\rho$, thereby enhancing the level of privacy protection.
@@ 6 Simulations
In this section, we conduct extensive simulations of a multi-robot system performing PP-ADRC both in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ to illustrate our theoretical results.
@@ 6.1 $2-$dimensional case
In this scenario, we deploy $10$ robots with a time-varying directed network, consisting of $2$ faulty robots and $8$ normal robots. The faulty robot sends its position to every normal robot. Each normal robot’s in-neighborhood is selected randomly at each iteration, while ensuring the conditions of resilience. The randomly generated initial states are depicted in Fig. 4(a), where blue points represent normal robots’ states and the red ones are the states of faulty robots. The green polygon represents the convex hull of normal robots. The faulty robots oscillate following the rule in [17]. For parameters, we set $\alpha_{i}(t)=0.8,\ \lambda=2.0$, and $\upsilon=0.75$.
@@
We execute the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 4(b), where blue points represent the final values each time, and the green point $(-0.011,0.210)$ is the sample expectation of blue points. We observe that the expectation falls in the convex hull of the initial states of normal robots.
Then, we only add noise to the second dimension to illustrate Theorem 4.17. We repeated the process $1000$ times, with $1000$ iterations for three different cases of noise parameters (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.5,\upsilon=0.85$) and obtain the convergence results shown in Fig. 5. The yellow convex hull represents $\mathcal{C}$, with $r_{2}=0.3$. The variances of the second dimension in order are $0.138,0.237,$ and $0.336$, respectively. In Fig. 5(d), the vertical axis represents the probability of the final value falling within the convex hull and the horizontal axis represents $r_{2}$. There are two sets of lines, i.e., one obtained through simulation probabilities $\mathrm{P}^{\rm Si}$ and the other $\mathrm{P}^{\rm Th}$ derived theoretically from Theorem 4.17. It is observed that in all three cases, the simulation results $\mathrm{P}^{\rm Si}$ are greater than the theoretical ones $\mathrm{P}^{\rm Th}$. When the distance satisifies $r_{2}=0.3$, we have $\delta(\mathcal{A},\mathcal{C})=0.919$ and $\mu(\mathcal{A})=1.77$, which illustrates our theoretical result $\delta(\mathcal{A},\mathcal{C})\leq\sqrt{\frac{d}{2}}\mu(\mathcal{A})+r_{2}$.
@@
At last, we illustrate the final value distribution using Mahalanobis distance. For the $2-$dimensional case, We also repeated the process $1000$ times, with $1000$ iterations for three different cases of noise parameters as below (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.0,\upsilon=0.65$), and obtain the convergence result shown in Fig. 6. Considering that the set of points equidistant by Mahalanobis distance forms an ellipse in two dimensions, we have drawn $4$ red ellipses under different parameters of $\chi$ for each figure, which are $2,3,4,5$ in order. It can be observed that $\mathrm{P}^{\mathrm{Si}}$ are greater than the theoretical one $\mathrm{P}^{\mathrm{Th}}$, illustrating our results in Theorem 4.6.
@@ 6.2 $3-$dimensional case
For the $3$-dimensional case, according to [34], although we can find a centerpoint with depth of $\frac{n}{d+1}=\frac{n}{4}$, but the running time is $O(n^{2})$, make the algorithm inefficient. Therefore, Har-Peled et al. proposed a compromised algorithm that achieves a depth of $\frac{n}{6}$ for the centerpoint with a runtime of $O(n\operatorname{log}n)$, which is used in our simulations here. We consider multi-robots rendezvous, where there are $12$ robots with a time-varying directed network, consisting of $2$ faulty robots and $10$ normal robots. Each normal robot interacts with one randomly chosen faulty robot and other normal robots, where the sufficient conditions for resilience are always guaranteed. The initial states are shown in Fig. 7(a), where the gray polyhedron represents the convex hull formed by normal robots. The faulty robots oscillate, with $\alpha_{i}(t)=0.8$, $\lambda=2.0$, and $\upsilon=0.75$, the same as the 2-dimensional case. We run the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 7(b). The blue points still represent final values each time, and the green point $(0.051,0.250,-0.312)$ is the sample expectation of final values. We can also observe that the expectation falls in the convex hull of the initial states of normal robots.
@@
Finally, we conduct the simulation to illustrate the final value distribution using Mahalanobis distance in the $3-$dimensional case shown in Fig. 9. The noise parameters are the same as the $2-$dimensional case, while for the gray hyperspheroid shown in the figure, we choose $\chi=3$. From Fig. 9(d), we observe that $\mathrm{P}^{\mathrm{Si}}$ are greater than the theoretical one $\mathrm{P}^{\mathrm{Th}}$, demonstrating out theoretical results in Theorem 4.6.

### v2

@@ 5.1 $\rho$-CGP in Resilient Vector Consensus
We can observe that the results do not contain any parameters related to faulty agents, indicating that their presence does not affect the privacy preservation outcome. While this may seem counterintuitive, it is actually reasonable. Each normal agent independently adds Gaussian noise to its local state before transmission to ensure privacy preservation, and this noise addition process is entirely unrelated to faulty agents. Once the information is transmitted, it may be intercepted by various entities, including faulty agents, honest-but-curious normal agents, and external eavesdroppers. However, all of these entities receive the same information—the noisy state transmitted by each normal agent—without any additional insights. Furthermore, the degree of privacy leakage of the initial states at nonzero iterations is determined by the step size $\gamma$ [24, 29]. Our analysis specifically addresses privacy preservation under these conditions, and thus the behavior of faulty agents does not impact the overall privacy preservation performance. Moreover, given the variables $\lambda$ and $\upsilon$, we plot the surface of $\rho$ concerning them, as shown in Fig. 3. It can be observed that $\rho$ monotonically decreases as both $\lambda$ and $\upsilon$ increase.
Remark 5.25.
Gaussian noise is canonical under CGP but may not be optimal. It is both interesting and essential to identify the optimal noise that best aligns with the specific requirements of the scenario, such as achieving the highest privacy preservation given the same noise strength. This problem presents significant challenges, especially in resilient vector consensus where we consider arbitrarily high-dimensional spaces. The complexity arises from two key aspects: First, the noise has numerous dimensional components, and second, these components can exhibit complicated coupling relationships across different dimensions. These characteristics substantially increase the difficulty in determining the optimal noise distribution. Addressing this challenging issue will be a major focus of our future research.
Figure 3: Influence of parameters $\upsilon$ and $\lambda$ on the coefficient $\rho$
5.2 Comparison with $(\varepsilon,\delta)$-DP
We first give the definition of $(\varepsilon,\delta)$-DP in resilient vector consensus.
@@ 5.2 Comparison with $(\varepsilon,\delta)$-DP
Given $\ell\in\mathbb{R}\geq 0$, the initial states of normal agents $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ are $\ell$-neighboring if, for some $i_{0}\in\overline{\mathcal{V}},||[\overline{\mathbf{x}}_{0}]_{i_{0}}-[\overline{\mathbf{x}}_{0}^{\prime}]_{i_{0}}||_{2}\leq\ell$ and $[\overline{\mathbf{x}}_{0}]_{i}=[\overline{\mathbf{x}}_{0}^{\prime}]_{i}$ if $i\neq i_{0}$ and $i\in\overline{\mathcal{V}}$. Given $\ell$, $\varepsilon$, and $\delta\in\mathbb{R}\geq 0$, the system satisfies $(\varepsilon,\delta)$-differential privacy if, for any pair of $\overline{\mathbf{x}}_{0}$ and $\overline{\mathbf{x}}_{0}^{\prime}$ of $\ell$-neighboring initial states and any set $\mathcal{O}\in\mathcal{B}((\mathbb{R}^{\overline{n}\times d})^{\mathbb{N}})$,
@@
The definition of neighboring initial states here differs from those in existing papers [25, 24, 29], where the $1$-norm is used. In those papers, differential privacy is $\varepsilon$-DP with Laplace noise, whereas we use $(\varepsilon,\delta)$-DP with Gaussian noise. Therefore, the neighboring initial states are defined using the $2$-norm.
As previously discussed, the system’s output is an infinite sequence. The output at each time can be considered as a query. Thus, the final result is the infinite composition of each single query at time $h$ satisfying $(\varepsilon(h),\delta(h))$-DP. We now detail $\varepsilon(h)$ and $\delta(h)$ according to the Theorem A.1. in [26].
@@
Regarding the composition, we can simply sum $\varepsilon(h)$ and $\delta(h)$ from $h=0$ to $\infty$. This leads to a rapid increase in privacy loss and $\delta$, resulting in a poor outcome. However, computing the tightest possible privacy guarantee for such a composition is #P-hard [44]. Then, we can make a detailed comparison between these two definitions below.
  • CGP better meets the privacy preservation needs. In resilient vector consensus, our privacy preservation target is the initial states of all the normal agents, which may be leaked during communication, rather than privacy leakage caused by the alteration or addition/deletion of an initial state in a data center. Therefore, the definition of $\ell$-neighboring states in DP does not fully meet our needs, as it should consider any pair of initial states. Consequently, CGP describes the degree of output differences for any pair of inputs, making it more suitable for protecting the initial state of each normal agent under resilient vector consensus. It also uses the distance between initial states as a privacy preservation parameter, thereby greatly improving the system’s utility.
@@ 5.3 Trade-off between Privacy and Accuracy
In this part, we discuss the trade-off between privacy and accuracy. As we analyze the convergence accuracy from two perspectives, i.e., distribution of the final value and convex hull change, we also detail the trade-off from these two perspectives. From Theorem 5.22, the parameters that determine $\rho$ are provided. The parameters that we can freely choose are the noise parameters $\lambda$ and $\upsilon$. The larger $\lambda$ and $\upsilon$ are, the stronger the privacy preservation is. Therefore, we analyze the trade-off between accuracy and privacy by varying $\lambda$ and $\upsilon$.
  • Distribution of the final value: The Mahalanobis distance is a normalized coefficient, meaning its value is fixed artificially and remains uninfluenced by variance or covariance. However, the volume of the hyperspheroid formed at an equal Mahalanobis distance can be influenced by various parameters. Therefore, the metric that accurately represents precision should be the volume of the hyperspheroid: the larger the volume, the lower the precision. As previously analyzed, increasing the amplitude parameters $\lambda$ and $\upsilon$ of the noise increases the volume of the hyperspheroid, thereby reducing convergence accuracy and decreasing $\rho$, which corresponds to enhanced privacy preservation.
  • Convex hull change: We use the Hausdorff distance to characterize the distance between convex hulls before and after adding noise. Similar to the Mahalanobis distance, the Hausdorff distance is also specified artificially. However, the probability of convergence within the convex hull is influenced by the noise parameters. The larger the noise amplitude parameters $\lambda$ and $\upsilon$, the lower the probability that the new convex hull will converge within a given Hausdorff distance. This results in a decrease in convergence accuracy but simultaneously leads to a reduction in $\rho$, thereby enhancing privacy preservation.
We can observe that although privacy and resilience are independent in algorithm design, their performance is tightly coupled. Identifying the appropriate noise parameters to achieve a balanced trade-off between them is crucial.
In fact, in addition to the trade-off between privacy and accuracy, there also exists a trade-off between privacy and convergence rate. From a qualitative perspective, increasing the noise magnitude enhances privacy but also slows convergence due to a slower decay of noise. However, we cannot establish a precise trade-off between privacy and convergence rate, as quantifying the convergence rate is infeasible. Analyzing the convergence rate presents two main challenges: First, while the centerpoint lies within the convex hull of the normal agents and is generally near the center, its exact parameters and convex combination bounds remain undetermined. This uncertainty complicates the calculation of agents’ movement at each iteration. As a result, most resilient vector consensus methods encounter these challenges and do not provide a quantitative analysis of the convergence rate, as in [2, 20, 18]. Second, the addition of noise increases complexity by introducing randomness into both the direction and magnitude of each agent’s movement, further complicating the analysis of the convergence rate.
@@ 6 Simulations
In this section, we consider the simulation on the real-world physical system, which is a multi-robot rendezvous [2]. We perform extensive simulations of a multi-robot system performing PP-ADRC in both $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ to illustrate our theoretical results.
@@ 6.1 $2$-dimensional case
In this scenario, we deploy $10$ robots with a time-varying directed network, consisting of $2$ faulty robots and $8$ normal robots. The messages sent by the two faulty agents to each normal agent are randomly selected within the region $[-0.7,-0.3]\times[0.3,0.7]$, causing the normal agents to converge to a point outside the convex hull, specifically at its upper-left corner. Each normal robot’s in-neighborhood is selected randomly at each iteration while ensuring the conditions of convergence. The randomly generated initial states are depicted in Fig. 4(a), where blue points represent normal robots’ states and the red ones are the states of faulty robots. The green polygon represents the convex hull of normal robots. For parameters, we set $\gamma_{i}(t)=0.8,\ \lambda=2.0$, and $\upsilon=0.75$.
@@
We execute the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 4(b), where blue points represent the final values each time, and the green point $(-0.160,0.239)$ is the sample mean of the blue points. We observe that the expectation falls in the convex hull of the initial states of the normal robots.
Then, we only add noise to the second dimension to illustrate Theorem 4.19. We repeat the process $1000$ times, with $1000$ iterations for three different cases of noise parameters (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.5,\upsilon=0.85$) and obtain the convergence results shown in Fig. 5. The yellow convex hull represents $\mathcal{C}$, with $r_{2}=0.3$. The variances of the second dimension in order are $0.066,0.117,$ and $0.219$, respectively. In Fig. 5(d), the vertical axis represents the probability of the final value falling within the convex hull and the horizontal axis represents $r_{2}$. There are two sets of lines, i.e., one obtained through simulation probabilities $\mathrm{P}^{\rm Si}$ and the other $\mathrm{P}^{\rm Th}$ derived theoretically from Theorem 4.19. It is observed that in all three cases, the simulation results $\mathrm{P}^{\rm Si}$ are greater than the theoretical ones $\mathrm{P}^{\rm Th}$. When the distance satisfies $r_{2}=0.3$, we have $\delta(\mathcal{A},\mathcal{C})=0.919$ and $\mu(\mathcal{A})=1.77$, which illustrates our theoretical result $\delta(\mathcal{A},\mathcal{C})\leq\sqrt{\frac{d}{2}}\mu(\mathcal{A})+r_{2}$.
@@
At last, we illustrate the final value distribution using Mahalanobis distance. We also repeat the process $1000$ times, with $1000$ iterations for three different cases of noise parameters (case 1): $\lambda=2.0,\upsilon=0.75$, case 2): $\lambda=2.5,\upsilon=0.75$, and case 3): $\lambda=2.0,\upsilon=0.65$), and obtain the convergence results shown in Fig. 6. Considering that the set of points equidistant in terms of Mahalanobis distance forms an ellipse in $\mathbb{R}^{2}$, we have drawn $4$ red ellipses under different parameters of $\chi$ for each figure, which are $2,3,4,5$ in order. There are also two sets of lines, i.e., one obtained through simulation probabilities $\mathrm{P}^{\rm Si}$ and the other $\mathrm{P}^{\rm Th}$ derived theoretically from Theorem 4.8. It is observed that in all three cases, the simulation results $\mathrm{P}^{\rm Si}$ are greater than the theoretical one $\mathrm{P}^{\rm Th}$.
@@ 6.2 $3$-dimensional case
For the $3$-dimensional case, according to [45], although we can find a centerpoint with depth of $\frac{n}{d+1}=\frac{n}{4}$, but the running time is $O(n^{2})$, make the algorithm inefficient. Therefore, Har-Peled et al. proposed a compromised algorithm that achieves a depth of $\frac{n}{6}$ for the centerpoint with a runtime of $O(n\operatorname{log}n)$, which is used in our simulations here. We consider multi-robot rendezvous, where there are $12$ robots with a time-varying directed network, consisting of $2$ faulty robots and $10$ normal robots. Each normal robot interacts with one randomly chosen faulty robot and other normal robots, where sufficient conditions for convergence are always guaranteed. The initial states are shown in Fig. 7(a), where the gray polyhedron represents the convex hull formed by the normal robots. We set $\gamma_{i}(t)=0.8$, $\lambda=2.0$, and $\upsilon=0.75$, consistent with the $2$-dimensional case. We run the algorithm $1000$ times, each with $1000$ iterations, and obtain the result in Fig. 7(b). The blue points represent the final values each time, and the green point $(0.051,0.250,-0.312)$ is the sample mean of the final values. We can observe that the expectation falls in the convex hull of the initial states of the normal robots.
@@
Finally, we conduct the simulation to illustrate the final value distribution using Mahalanobis distance in the $3$-dimensional case shown in Fig. 9. The noise parameters are identical to those in the $2$-dimensional case. For the gray hyperspheroids shown in the figure, we choose $\chi=3$. From Fig. 9(d), we observe that $\mathrm{P}^{\mathrm{Si}}$ are greater than the theoretical one $\mathrm{P}^{\mathrm{Th}}$, demonstrating out theoretical results in Theorem 4.8.
@@
6.3 Privacy preserving perfomance
We consider a system with five normal robots and one faulty robot in a complete graph. The faulty robot sends arbitrary, different messages to each normal robot, randomly chosen from $[-0.7,-0.4]\times[0.4,0.7]$. Each normal robot exchanges state information with all others. The parameters are set as $\lambda=2.0$, $\upsilon=0.75$, and $\gamma_{l}=0.4$. The algorithm is executed $1000$ times, each with $1000$ iterations, for two initial states shown in Fig. 10(a). The initial state distance is $\mathrm{dist}(\overline{x}_{0},\overline{x}_{0}^{\prime})=0.2236$. Normal robots in group $1$ start at $\overline{x}_{0}$, while those in group $2$ start at $\overline{x}_{0}^{\prime}$. The results are presented in Fig. 10(b). Blue circles indicate the final values of group $1$, with the green circle as their sample mean. Blue crosses denote the final values of group $2$, and the green triangle represents their sample mean.
(a) Comparison of $\overline{x}_{0}$ and $\overline{x}_{0}^{\prime}$
(b) Comparison of final values
Figure 10: Simulations for initial states $\overline{x}_{0}$ and $\overline{x}_{0}^{\prime}$
Substituting the parameters into the result of Theorem 5.22, we obtain the following expression for the Rényi divergence.
$D_{\alpha}(\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\|\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}))\leq 0.9374\alpha.$ (28)
We collect all the data transmitted by the normal robots over $1000$ iterations and repeat this process $1000$ times to fit the distribution. Finally, by setting $\alpha=2.0$, we obtain a Rényi divergence of $D_{\alpha}(\mathbf{y}_{\overline{\mathbf{x}}_{0}}(\mathbf{N})\|\mathbf{y}_{\overline{\mathbf{x}}_{0}^{\prime}}(\mathbf{N}))=1.2863$ between the two distributions, which is smaller than $0.9374\times 2=1.8748$.